OpenStudy (anonymous):

It takes an electrician 45 minutes to wire an office in a new office building that is being built. The electrician signed a contract to wire all offices on 6 floors of the new building. There are 50 offices on each floor. How many weeks should the electrician allow for this job if he works a standard work week (8 hours per day for 5 days a week)? Express the answer in decimal form rounded to the tenths place

OpenStudy (anonymous):

This looks like a problem with a lot of steps... Try starting by finding how long it will take to wire all 50 offices on one floor if it takes 45 minutes for each office.

OpenStudy (anonymous):

So maybe it would look something like this: 45 minutes to wire 1 office... 45 minutes is 45/60 = 3/4 of an hour, or 0.75 hours. so:\[\frac{ 0.75 }{ 1 }\frac{ hour }{ office } = \frac{ ? }{ 50 }\frac{ hours }{ offices }\]

OpenStudy (anonymous):

The answer is going to be that it is going to take him 90 days to complete the job. So divide that by 7.... and then that will tell you how many weeks.